Accelerator Update for December 19, 2007 - December 21, 2007
Wednesday December 19
The day shift began with Operations monitoring the Tevatron (TeV) store 5809 (37.35E30), the Recycler stash (122.1E10), and the Antiproton Source (Pbar) stack (190.2E10), with Pbar conducting studies, with the Main Injector (MI), Booster, and the Linac operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams, and with NuMI , SciBooNE / MiniBooNE experiments taking beam.
At 9:33 AM, CDF and D0 turned off their detectors for a TeV end-of-store study.
At 9:47 AM, LRF2 tripped off. It reset.
Pbar experts halted their study at 11:05 AM.
Linac experts reported at 11:16 AM that they had completed repairs on a quadrupole power supply (QPS107) and on the KRF2 flow sensor. Operations resumed sending beam to Pbar, Sci/MiniBooNE, NuMI, and SY.
Operations terminated store 5809 at 12:08 PM.
Operations began preparing D0 for access at 12:19 PM.
Operations prepared the Meson M03 & 4 enclosures for access at 1:48 PM.
Pbar experts resumed their studies at 2:09 PM.
Operations resumed sending beam to Meson at 2:26 PM.
Operations prepared M03 & 4 for access at 2:59 PM.
Operations reported at 3:24 PM that SY experts had completed the special MTest run.
Operations put the TeV through a dry squeeze at 3:33 PM.
Pbar experts completed their studies at 3:55 PM.
Operations reported at 4:11 PM that the Recycler had stashed antiprotons twice during the day shift.
Operations put the TeV through a wet squeeze at 4:31 PM.
At 4:51 PM, Operations reported that the L:QPS107 power supply was wandering. An operator adjusted it back into tolerance.
Operations put the TeV into shot setup at 5:18 PM.
At 6:49 PM, Operations reported that the TeV had aborted five times due to bad beam positions at TeV sector E4. Operations contacted an expert. The problems were due to the limits set for the E4 BLMs, which were adjusted.
Operations established store 5811 at 8:34 PM with an initial luminosity of 191.54E30.
At 10:13 PM, the Pbar lithium lens tripped off. It reset.
At 10:40 PM, QPS107 began wandering again.
At 11:25 PM, the Pbar lithium lens tripped off. It reset.
Operations reported at midnight that the Recycler had stashed twice during the evening shift.
Thursday December 20
The midnight shift began with the TeV recovering from the C4 quench, with Operations monitoring the stash (98.0E10), with Pbar stacking (32.9E10), and with NuMI and Sci /MiniBooNE taking beam.
Operations reported at 1:15 AM that the TeV sector C4 cold compressor had started running erratically. The wet engine tried to keep up, but the rest of the cryo system couldn’t recover. Cryo experts investigated.
At 1:16 AM, the Pbar lithium lens tripped off. It reset.
At 1:24 AM, the L:QPS107 power supply wandered again. An operator adjusted it.
At 1:33 AM, the Pbar lithium lens tripped off for the first of two times in 20 minutes. It would not reset after the last trip. Operations called in an expert.
At 2:07 AM, LRF3 tripped off. It reset.
A Pbar expert reported at 3:35 AM that he had replaced a failed meter. Pbar resumed stacking a few minutes later.
At 3:58 AM, KRF4 tripped off. It reset.
At 4:26 AM, the L:QPS107 power supply drifted out of tolerance again. An operator adjusted.
Cryo technicians reported at 7:29 AM that the C4 cryo system was back online.
Operations reported at 8:05 AM that the Recycler had stashed four times during the midnight shift.
Operations put the TeV through a dry squeeze at 8:40 AM.
At 8:53 AM, Operations reported that two Linac quadrupoles (QPS107 & 511) had been causing bad beam to Booster.
Operations put the TeV into shot setup at 9:50 AM.
The Neutron Therapy Facility (NTF) reported at 10:52 AM that they had finished treating their last patient for the day.
Operations established store 5813 at 12:30 PM with an initial luminosity of 183.23E30.
SY experts began taking beam at 1:54 PM and tuned the beam to the MTest area.
At 2:44 PM, BRF5 tripped off and would not reset. Operators bypassed the station and brought BRF19 on in its place.
Operations reported at 4:12 PM that the Recycler had stashed two times during the day shift.
The MI slipstacking and coalescing studies ended at 4:58 PM.
The MI-30 collimator study ended at 5:29 PM.
SY experts completed their MTest beam tuning at 5:37 PM.
At 7:45 PM, a Booster power supply (B:MH2) tripped off for the first of three times in about 30 minutes. There was never an indication of what was causing the trips. It was reset after each trip.
At 8:41 PM, KRF5 tripped off for the first of two times in about 25 minutes. It tripped off on a water flow indication. It reset.
At 8:48 PM, B:MH2 tripped. An expert asked Operations to check the PLC status inside the supply. There was one bad bit: excessive ripple. Operations contacted the expert.
At 9:31 PM, Operations reported that L:QPS107 had gone out of tolerance 90 times in that last eight hours.
At 10:47 PM, B:MH2 tripped off. A Booster expert was here and he investigated.
At 11:17 PM, a TeV sector E48 Kicker tripped off. Operations contacted an expert.
At 11:37 PM, the Pbar lithium lens tripped off. It reset.
At 11:40 PM, B:MH2 tripped off again. The Booster experts replaced a cable driver card and reset the supply.
Operations reported at midnight that the Recycler had stashed four times during the evening shift.
Friday December 21
The midnight shift began with Operations monitoring store 5813 (51.86E30) and a stash (200.6E10), with Pbar stacking (11.2E10), and with NuMI and Sci/MiniBooNE taking beam.
At 12:27 AM, B:MH2 tripped off. It reset.
At 12:30 AM, KRF4 tripped off. It reset.
At 12:34 AM, a power glitch caused the Booster West Anode, KRF3, 4 & 7, the Pbar lens, TEL, and a TeV sector A4 roughing pump to trip.
Linac and Booster resumed sending beam at 12:38 AM.
Operators had the TEL back on by 12:44 AM.
At 2:48 AM, the cryo system at A4 went into oscillation, but it seemed to be holding its own.
At 2:51 AM, the TeV quenched. Vacuum had been stable, but then went bad very quickly. Operators turned the A4 vacuum pump back on.
At 3:58 AM, Operations reported that the vacuum at A4 was not improving. Operations called in vacuum experts to add more pumps.
Operations prepared the TeV A-E areas for access at 4 AM.
TeV experts accessed at 6:21 AM to change out the surge resistor on the C17 separator.
At 7:25 AM, BRF10 tripped off. It reset.
At 7:53 AM, LRF5 tripped off. It reset.
At 7:55 AM, Operations reported that L:QPS107 had gone out of tolerance 109 times since 9:30 PM on Thursday night.
Operations reported at 7:56 AM that the Recycler had stashed four times during the midnight shift.
The Plans for Friday and the Weekend
The plans for today are to fix the TeV sector A4 vacuum, the Linac quadrupole power supply (QPS107), and then resume stacking, stashing, and storing for today and the weekend.
Comments and Suggestions
|last modified 12/21/2007 email Fermilab|
|Security, Privacy, Legal|